# A Study On The Graph Formulation Of Union Closed Conjecture

### v2

@@ 3 Some New Graph Results
We will determine the number of maximal stable sets using Lemma 3.5. First, let’s apply Lemma 3.3 to Equation 5.
$w_{\scriptscriptstyle G}$ $=\sum_{\scriptscriptstyle\Theta\subseteq[n]}\sum_{\scriptscriptstyle\Gamma\subseteq\Theta^{\mathsf{c}}}\left\{w_{\scriptscriptstyle C\setminus\Gamma}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma)})\sum_{\scriptscriptstyle\Psi\subseteq\Theta^{\mathsf{c}}\setminus\Gamma}w_{\scriptscriptstyle H\setminus\Psi}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Psi)})\right\}$ (8)
$=\sum_{\scriptscriptstyle\Theta\subseteq[n]}\sum_{\scriptscriptstyle\Gamma\subseteq\Theta^{\mathsf{c}}}\left\{w_{\scriptscriptstyle C}(\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})\sum_{\scriptscriptstyle\Psi\subseteq\Theta^{\mathsf{c}}\setminus\Gamma}w_{\scriptscriptstyle H}(\Theta\cup\Psi,\overline{\Theta^{\mathsf{c}}\setminus\Psi})\right\}$ (9)
‘ Notice that,
@@
Observe that this identity holds because we are summing over all possible subsets $\Psi\subseteq\Theta^{\mathsf{c}}\setminus\Gamma$. Next, we introduce a new variable $\Omega=\Theta\cup\Gamma$, with $\Omega^{\mathsf{c}}=[n]\setminus\Omega=\Theta^{\mathsf{c}}\setminus\Gamma$. This allows us to express $w_{\scriptscriptstyle C}(\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})$ as $w_{\scriptscriptstyle C}(\Omega,\overline{\Omega^{\mathsf{c}}})$. We can now apply this change of variables after substituting (11) into equation (9).
@@
$w_{\scriptscriptstyle G}$ $=\sum_{\scriptscriptstyle\Theta\subseteq[n]}\sum_{\scriptscriptstyle\Gamma\subseteq\Theta^{\mathsf{c}}}\left\{w_{\scriptscriptstyle C}(\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})w_{\scriptscriptstyle H}(\Theta,\overline{\Gamma})\right\}$ ([Substituting equation 11])
@@
$=\sum_{\scriptscriptstyle\Omega\subseteq[n]}\sum_{\scriptscriptstyle\Theta\subseteq\Omega}\left\{w_{\scriptscriptstyle C}(\Omega,\overline{\Omega^{\mathsf{c}}})w_{\scriptscriptstyle H}(\Theta,\overline{\Omega\setminus\Theta})\right\}$ ([Change of variables])
@@
$=\sum_{\scriptscriptstyle\Omega\subseteq[n]}\left\{w_{\scriptscriptstyle C}(\Omega,\overline{\Omega^{\mathsf{c}}})\sum_{\scriptscriptstyle\Theta\subseteq\Omega}w_{\scriptscriptstyle H}(\Theta,\overline{\Omega\setminus\Theta})\right\}$
$=\sum_{\scriptscriptstyle\Omega\subseteq[n]}\left\{w_{\scriptscriptstyle C}(\Omega,\overline{\Omega^{\mathsf{c}}})w_{\scriptscriptstyle H}\right\}$
$=w_{\scriptscriptstyle C}w_{\scriptscriptstyle H}$ (12)
In a similar manner, it can be deduced that $w_{\scriptscriptstyle G}(b)=w_{\scriptscriptstyle C}(b)w_{\scriptscriptstyle H}$ for any $b\in C\setminus[n]$. When $a\in[n]$, for every $\Gamma$ with $a\in\Gamma$, we have $w_{\scriptscriptstyle C\setminus\Gamma}(\{a\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle C}(\Gamma)})=0$. Using this fact and after similar calculations, one can find that $w_{\scriptscriptstyle G}(a)\leq w_{\scriptscriptstyle C}(a)w_{\scriptscriptstyle H}(a)$. Let $r\in C$ be a rare vertex in $C$. If $r\in C\setminus[n]$, we have $w_{\scriptscriptstyle G}-2w_{\scriptscriptstyle G}(r)=w_{\scriptscriptstyle C}w_{\scriptscriptstyle H}-2w_{\scriptscriptstyle C}(r)w_{\scriptscriptstyle H}=(w_{\scriptscriptstyle C}-2w_{\scriptscriptstyle C}(r))w_{\scriptscriptstyle H}\geq 0$. Now, if $r\in[n]$, then $w_{\scriptscriptstyle G}-2w_{\scriptscriptstyle G}(r)\geq w_{\scriptscriptstyle C}w_{\scriptscriptstyle H}-2w_{\scriptscriptstyle C}(r)w_{\scriptscriptstyle H}(r)\geq(w_{\scriptscriptstyle C}-2w_{\scriptscriptstyle C}(r))w_{\scriptscriptstyle H}\geq 0$. Therefore, every rare vertex in $C$ is also rare in $G$. ∎
Figure 1: Three merged graphs, with a 2-layered vertex highlighted in red. The resultant graph satisfies the conjecture because each of the three individual graphs satisfies the conjecture. See Proposition 3.11.
@@
Figure 2: The bipartite graph shown consists of two bipartite classes represented by red and blue vertices. As all the vertices in red bipartite class has a pendant vertex adjacent to it, the graph satisfies the conjecture. See Corollary 3.9.
@@
The theorem’s implications extend to solving the conjecture for previously unexplored graph classes, opening up new avenues of research. Now, let’s discuss some of the notable consequences stemming from this result. All results concerning the set version of the union closed conjecture presented below are directly derived from the corresponding graph results by applying Proposition 2.1 within the respective graph results.
@@
Let $\mathcal{F}_{\scriptscriptstyle 1}$ and $\mathcal{F}_{\scriptscriptstyle 2}$ be two families of sets that do not share any common sets. Let $A$ denote the set of elements that are common to both families, i.e., $A=U(\mathcal{F}_{\scriptscriptstyle 1})\cap U(\mathcal{F}_{\scriptscriptstyle 2})$. Assume that every $S\in\mathcal{F}_{\scriptscriptstyle 1}\cup\mathcal{F}_{\scriptscriptstyle 2}$ with $S\cap A\neq\emptyset$ contains an element of frequency 1. If $\langle\mathcal{F}_{\scriptscriptstyle 1}\rangle$ or $\langle\mathcal{F}_{\scriptscriptstyle 2}\rangle$ satisfies the union closed conjecture, then $\langle\mathcal{F}_{\scriptscriptstyle 1}\cup\mathcal{F}_{\scriptscriptstyle 2}\rangle$ also satisfies the conjecture.
@@
Let $G$ be a bipartite graph and $v\in V(G)$ be such that there exists at least one pendent vertex adjacent to $v$ and all the non-pendent neighbors of $v$ are 2-layered. Then $G$ satisfies Frankl’s conjecture. Furthermore, the vertex $v$ and all of its neighbors are rare in $G$.
@@
Define $P=\{u\mid u\in N_{\scriptscriptstyle G}(v),d(u)=1\}$, where $d(u)$ denotes the degree of vertex $u$. Let $C$ be the subgraph induced by vertex $v$, while $H$ is the subgraph induced by the vertex subset $V(G)\setminus(P\cup\{v\})$. Clearly, $V(H)\cap V(C)=N_{\scriptscriptstyle G}(v)\setminus P$, which corresponds to the non-pendant neighbors of $v$.
Observe that all vertices in $N(v)\setminus P$ are 2-layered vertices, and $\{C,H\}$ forms a decomposition of $G$. It is important to note that $C$ is a star graph, implying that all of its vertices are rare within $C$. Consequently, by applying Theorem 3.2, we can conclude that both vertex $v$ and its neighbors in $G$ are rare in $G$, thereby establishing that $G$ satisfies the conjecture. ∎
@@
Let $\mathcal{F}$ be a family of sets. Consider a member set $A\in\mathcal{F}$ that contains at least one element with frequency one. Suppose that for every set $S\in\mathcal{F}$, where $S\cap A\neq\emptyset$, there exists an element within $S$ that has frequency one. Under these conditions, the family $\langle\mathcal{F}\rangle$ satisfies the union closed conjecture, with all elements of $A$ being abundant.
@@
Let $G$ be a bipartite graph with no isolated vertices, in which every vertex in one of its bipartite classes is adjacent to at least one pendant vertex. Then, $G$ satisfies Frankl’s conjecture. In fact, all the vertices of $G$ are rare.
@@
Let $\mathcal{F}$ be a family of sets where each of it’s member sets containing an element of frequency 1. Then $\langle\mathcal{F}\rangle$ satisfies the union closed conjecture.
@@ 4 Acknowledgments
I would like to express my sincere gratitude to my adviser, Prof. Rogers Mathew, for his invaluable guidance.

### v4

@@ 3 Some New Graph Results
Let $m\in V(H)\cap V(I)$, and suppose $N_{\scriptscriptstyle I}(m)=\emptyset$. In this case, we can simply remove the vertex from $I$ to get a new decomposition. Similarly, if $N_{\scriptscriptstyle H}(m)=\emptyset$, we can remove $m$ from $H$. This allows us to redefine $H$ and $I$ in such a way that every vertex in $V(H)\cap V(I)$ has neighbors in both $H$ and $I$. Let $[n]$ be the vertices common to these subgraphs $H$ and $I$. We will determine the number of maximal stable sets using Lemma 3.6. First, let’s apply Lemma 3.4 to Equation 5.
$w_{\scriptscriptstyle G}$ $=\sum_{\scriptscriptstyle\Theta\subseteq[n]}\sum_{\scriptscriptstyle\Gamma\subseteq\Theta^{\mathsf{c}}}\left\{w_{\scriptscriptstyle I\setminus\Gamma}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle I}(\Gamma)})\sum_{\scriptscriptstyle\Psi\subseteq\Theta^{\mathsf{c}}\setminus\Gamma}w_{\scriptscriptstyle H\setminus\Psi}(\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle H}(\Psi)})\right\}$ (8)
$=\sum_{\scriptscriptstyle\Theta\subseteq[n]}\sum_{\scriptscriptstyle\Gamma\subseteq\Theta^{\mathsf{c}}}\left\{w_{\scriptscriptstyle I}(\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})\sum_{\scriptscriptstyle\Psi\subseteq\Theta^{\mathsf{c}}\setminus\Gamma}w_{\scriptscriptstyle H}(\Theta\cup\Psi,\overline{\Theta^{\mathsf{c}}\setminus\Psi})\right\}$ (9)
Notice that,
@@
This identity holds because we are summing over all possible subsets $\Psi\subseteq\Theta^{\mathsf{c}}\setminus\Gamma$. Next, we introduce a new variable $\Omega=\Theta\cup\Gamma$, with $\Omega^{\mathsf{c}}=[n]\setminus\Omega=\Theta^{\mathsf{c}}\setminus\Gamma$. This allows us to express $w_{\scriptscriptstyle I}(\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})$ as $w_{\scriptscriptstyle I}(\Omega,\overline{\Omega^{\mathsf{c}}})$. We can now apply this change of variables after substituting (11) into equation (9).
@@
$w_{\scriptscriptstyle G}$ $=\sum_{\scriptscriptstyle\Theta\subseteq[n]}\sum_{\scriptscriptstyle\Gamma\subseteq\Theta^{\mathsf{c}}}\left\{w_{\scriptscriptstyle I}(\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})w_{\scriptscriptstyle H}(\Theta,\overline{\Gamma})\right\}$ ([Substituting equation 11])
@@
$=\sum_{\scriptscriptstyle\Omega\subseteq[n]}\sum_{\scriptscriptstyle\Theta\subseteq\Omega}\left\{w_{\scriptscriptstyle I}(\Omega,\overline{\Omega^{\mathsf{c}}})w_{\scriptscriptstyle H}(\Theta,\overline{\Omega\setminus\Theta})\right\}$ ([Change of variables])
@@
$=\sum_{\scriptscriptstyle\Omega\subseteq[n]}\left\{w_{\scriptscriptstyle I}(\Omega,\overline{\Omega^{\mathsf{c}}})\sum_{\scriptscriptstyle\Theta\subseteq\Omega}w_{\scriptscriptstyle H}(\Theta,\overline{\Omega\setminus\Theta})\right\}$
$=\sum_{\scriptscriptstyle\Omega\subseteq[n]}\left\{w_{\scriptscriptstyle I}(\Omega,\overline{\Omega^{\mathsf{c}}})w_{\scriptscriptstyle H}\right\}$
$=w_{\scriptscriptstyle I}w_{\scriptscriptstyle H}$ (12)
Lemma 3.7.
Let $G$ be a graph with a decomposition $\{H,I\}$ such that $V(H)\cap V(I)=[n]$, where $[n]$ is a stable set. Assume that no vertex of $[n]$ is isolated in either $H$ or $I$, and that every vertex of $[n]$ is 2-layered. Let $a\in V(I)\setminus[n]$ and $b\in[n]$. Then,
$w_{\scriptscriptstyle G}(a)=w_{\scriptscriptstyle I}(a)w_{\scriptscriptstyle H}$ (13)
$w_{\scriptscriptstyle G}(b)\leq w_{\scriptscriptstyle I}(b)w_{\scriptscriptstyle H}(b)$ (14)
Proof.
Very similar to the proof of Theorem 3.2, applying Lemma 3.4 to Equation 6, it can be deduced that $w_{\scriptscriptstyle G}(a)=w_{\scriptscriptstyle I}(a)\,w_{\scriptscriptstyle H}$ for any $a\in V(I)\setminus[n]$. In order to prove Equation 14, we apply Lemma 3.4 to Equation 7. Now consider a vertex $b\in[n]$. One should note that Lemma 3.4 states $w_{\scriptscriptstyle I\setminus\Gamma}(\{b\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle I}(\Gamma)})=w_{\scriptscriptstyle I}(\{b\}\cup\Theta\cup\Gamma,\overline{\Theta^{\mathsf{c}}\setminus\Gamma})$ whenever $b\notin\Gamma$. Observe that for every $\Gamma$ with $b\in\Gamma$, we have $w_{\scriptscriptstyle I\setminus\Gamma}(\{b\}\cup\Theta,\overline{\Theta^{\mathsf{c}}\cup N_{\scriptscriptstyle I}(\Gamma)})=0$. Therefore, some terms vanish, and since the steps are otherwise identical to the proof of Theorem 3.2, we obtain $w_{\scriptscriptstyle G}(b)\leq w_{\scriptscriptstyle I}(b)\,w_{\scriptscriptstyle H}(b)$. ∎
Proof of Theorem 3.3.
We redefine $H$ and $I$ so that no common vertex is isolated in either subgraph, as in the proof of Theorem 3.2. Let $[n]$ denote the set of common vertices. Since all vertices of $[n]$ belong to the same bipartite class, $[n]$ forms a stable set. Moreover, as each vertex of $[n]$ is 2-layered, we may apply Theorem 3.2 and Lemma 3.7. Consider $r\in V(I)$ a rare vertex in $I$. If $r\in V(I)\setminus[n]$, we have $w_{\scriptscriptstyle G}-2w_{\scriptscriptstyle G}(r)=w_{\scriptscriptstyle I}w_{\scriptscriptstyle H}-2w_{\scriptscriptstyle I}(r)w_{\scriptscriptstyle H}=(w_{\scriptscriptstyle I}-2w_{\scriptscriptstyle I}(r))w_{\scriptscriptstyle H}\geq 0$. Now, if $r\in[n]$, then $w_{\scriptscriptstyle G}-2w_{\scriptscriptstyle G}(r)\geq w_{\scriptscriptstyle I}w_{\scriptscriptstyle H}-2w_{\scriptscriptstyle I}(r)w_{\scriptscriptstyle H}(r)\geq(w_{\scriptscriptstyle I}-2w_{\scriptscriptstyle I}(r))w_{\scriptscriptstyle H}\geq 0$. Therefore, every rare vertex in $I$ is also rare in $G$. ∎
@@
Figure 5: The bipartite graph shown consists of two bipartite classes represented by red and blue vertices. As all the vertices in red bipartite class has a pendant vertex adjacent to it, the graph satisfies the conjecture.
Figure 6: Three merged graphs, with a 2-layered vertex highlighted in red. The resultant graph satisfies the conjecture because each of the three individual graphs satisfies the conjecture.
@@
The theorem’s implications extend to solving the conjecture for previously unexplored graph classes, opening up new avenues of research. Now, let’s discuss some of the notable consequences stemming from this result. All results concerning the set version of the Union Closed Sets Conjecture presented below are directly derived from the corresponding graph results by applying Proposition 2.1 within the respective graph results.
@@
Let $\mathcal{F}_{\scriptscriptstyle 1}$ and $\mathcal{F}_{\scriptscriptstyle 2}$ be two families of sets that do not share any common sets. Let $A$ denote the set of elements that are common to both families, i.e., $A=U(\mathcal{F}_{\scriptscriptstyle 1})\cap U(\mathcal{F}_{\scriptscriptstyle 2})$. Assume that every $S\in\mathcal{F}_{\scriptscriptstyle 1}\cup\mathcal{F}_{\scriptscriptstyle 2}$ with $S\cap A\neq\emptyset$ contains an element of frequency 1. If $\langle\mathcal{F}_{\scriptscriptstyle 1}\rangle$ or $\langle\mathcal{F}_{\scriptscriptstyle 2}\rangle$ satisfies the UCC condition, then $\langle\mathcal{F}_{\scriptscriptstyle 1}\cup\mathcal{F}_{\scriptscriptstyle 2}\rangle$ also satisfies the UCC condition.
@@
Let $G$ be a bipartite graph and $v\in V(G)$ be such that there exists at least one pendant vertex adjacent to $v$ and all the non-pendant neighbors of $v$ are 2-layered. Then $G$ satisfies UCC condition. Furthermore, the vertex $v$ and all of its neighbors are rare in $G$.
@@
Define the set of pendant neighbors of $v$ as $P=\{u\in N_{\scriptscriptstyle G}(v)\mid d(u)=1\},$where $d(u)$ denotes the degree of $u$. Let $I$ be the subgraph induced by vertex $v$, while $H$ is the subgraph induced by the vertex subset $V(G)\setminus(P\cup\{v\})$. Clearly, $V(H)\cap V(I)=N_{\scriptscriptstyle G}(v)\setminus P$, which corresponds to the non-pendant neighbors of $v$.
Observe that all vertices in $N(v)\setminus P$ are 2-layered vertices, and $\{I,H\}$ forms a decomposition of $G$. It is important to note that $I$ is a star graph, implying that all of its vertices are rare within $I$. Consequently, by applying Theorem 3.3, we can conclude that both vertex $v$ and its neighbors in $G$ are rare in $G$, thereby establishing that $G$ satisfies the conjecture. ∎
@@
Let $\mathcal{F}$ be a family of sets. Consider a member set $A\in\mathcal{F}$ that contains at least one element with frequency one. Suppose that for every set $S\in\mathcal{F}$, where $S\cap A\neq\emptyset$, there exists an element within $S$ that has frequency one. Under these conditions, the family $\langle\mathcal{F}\rangle$ satisfies the UCC condition, with all elements of $A$ being abundant.
@@
Let $G$ be a bipartite graph with no isolated vertices, in which every vertex in one of its bipartite classes is adjacent to at least one pendant vertex. Then, $G$ satisfies UCC condition. In fact, all the vertices of $G$ are rare.
@@
Let $\mathcal{F}$ be a family of sets where each of its member sets contains an element of frequency 1. Then $\langle\mathcal{F}\rangle$ satisfies the UCC condition.
@@
As an illustration, Figure 6 provides an example of Corollary 3.11, while Figure 6 demonstrates an example of Proposition 3.13.
@@
4 Future Research
Several graph classes, such as subcubic graphs, chordal bipartite graphs, and series-parallel graphs, are already known to satisfy the Union Closed Sets Conjecture. However, many families still remain unresolved; for example, the conjecture has not yet been established even for the classes of regular bipartite graphs, planar bipartite graphs, and bipartite graphs of treewidth three. This represents an important direction for further exploration.
In this paper, we identified some additional graph classes that satisfy the conjecture, but more significantly, we developed a proof technique that may inspire further work. Our method crucially requires that the common vertices in the decomposition be 2-layered. A natural next step is to investigate situations where some, or even all, of these vertices fail to be 2-layered. For instance, one may ask whether there exist particular cases in which the number of maximal stable sets equals the product of the numbers of maximal stable sets of the decompositions, despite the lack of the 2-layered property. Establishing such results could pave the way toward a more general theorem.
Another promising perspective is to study the conjecture by systematically decomposing graphs into two or more subgraphs and analyzing the structure of maximal stable sets within each piece. Even in cases where the common vertices are not 2-layered, one can investigate whether useful bounds can still be obtained using ideas similar to those presented here. To the best of our knowledge, the approach of analyzing the Union Closed Sets Conjecture through graph decompositions has not been explored in the literature. Therefore, pursuing this direction may open new avenues for understanding the conjecture.
@@ 5 Acknowledgments
I would like to express my sincere gratitude to my adviser, Prof. Rogers Mathew, for his invaluable guidance and support.